# E2CL: Exploration-based Error Correction Learning for Embodied Agents

### v1

@@ 3.3 Ablation Study on Training Data
In this section, we explore the impact of the collected training data, i.e., feedback data $D_{f}$ and correction data $D_{c}$, on overall performance by ablating them in the training. During the inference phase, we employ speculative inference for all settings to ensure consistency. As shown in Table 2, we observe that both $D_{f}$ and $D_{c}$ are each beneficial for the agent, but lag behind the combination of them. We hypothesize that the improvement observed when training with $D_{c}$ is primarily due to the enhanced self-correction capability of the agent. However, the limited ability to generate high-quality action feedback hampers the effectiveness of self-correction during speculative inference, as demonstrated in Section 3.6. Compared to the agent training without both $D_{f}$ and $D_{c}$, training with $D_{p}$ and $D_{f}$ improves the performance by training to predict environmental feedback, which explicitly aligns with environmental knowledge. However, the weak self-correction capability of the agent constrains the agent from generating executable and correct action in speculative inference, which is demonstrated in  Section 3.5. In our method, we integrate both types of data, enabling our agent to generate higher-quality action feedback and exhibit stronger self-correction abilities. This results in a substantial performance boost compared to other ablation settings.
@@ 3.4 Analysis on Different Size of the Model
To investigate the impact of model size on performance, we train models of different sizes on seen tasks using both BC and our method, and evaluated them on unseen tasks. Our results are in line with common experience, where larger models perform relatively better across all aspects, indicating that model scale significantly impacts performance. Moreover, it can also be observed that our method outperforms BC in both Affordance rate and LCS across models with different parameter sizes, which demonstrates that our method consistently provides superior performance regardless of model size. Notably, when using our method, smaller models achieve performance surpassing larger models using BC across all metrics. This finding suggests that our method is able to release the potential of small language models and lays the foundation for building agents that work on edge devices in the future.
@@ 3.5 Evaluation on Self-Correction Ability
3.6 Analysis on Speculative Inference
To analyze the contribution of speculative inference to overall performance, as well as to explore the quality and effectiveness of self-generated feedback, we conduct an analysis on speculative inference.
@@ 3.6 Analysis on Speculative Inference
Table 3: Performance of the agent without speculative inference. "si" denotes speculative inference as described.
@@ 3.7 Error Analysis
We also perform an error analysis to identify the aspects where the agent constructed using our method outperforms BC. There are a total of eight types of errors, which can be further classified into grounding errors (object availability) and execution-related errors (others). The detailed demonstration can be found in the Appendix B. As shown in Figure 6, we observe that all error types decreased by more than 24%, with Over occupied error showing the highest reduction rate of 94.4%. This highlights the effectiveness of our method in reducing various types of errors, highlighting its comprehensiveness. For the two most frequent types of execution-related errors, unflipped boolean state and agent proximity, our method achieves a reduction in error count by over 37% compared to BC, thereby demonstrating its effectiveness. Although our method primarily aims to avoid execution errors related to physical constraint and does not specifically target grounding errors such as object availability, the fact that it still reduces this type of error demonstrates the generalizability of our method.
@@ 4 Related Work
After exploration, the agent would encounter failure in the past experience, which is assumed as negative samples. The topic of learning from negative samples has increasingly gained attention as an alternative approach to learning solely from positive samples. Traditionally, some studies aim to decrease the probability of negative samples while increasing the probability of positive samples in order to achieve better performance (Wang et al., 2024a; Zheng et al., 2023; Liu et al., 2022). Additionally, some works construct correcting dataset and tuning language models on these data (An et al., 2023; Wang et al., 2024b; Bai et al., 2022). Besides, there are other efforts aimed at leveraging the comprehension abilities of language models to widen the gap between positive and negative samples (An et al., 2023; Liu et al., 2023; Zhang et al., 2023). In our work, we similarly leverage the inherent understanding capabilities of language models and enhance the embodied agent’s learning from environmental feedback on exploration errors, as well as its ability to self-correct.
@@ 5 Conclusion
In this work, we aim to align the embodied agent with the environment to enhance its task-solving performance. Firstly, we present E2CL, a novel framework that leverages exploration-induced errors and environmental feedback to enhance environment alignment for LM-based agents during teacher-guided and teacher-free exploration. Furthermore, we introduce speculative inference, a process in which the agent utilizes learned abilities for self-feedback and self-correction to reduce execution errors. Extensive experiments show that our method outperforms behavior cloning and other baseline methods.
@@
The baseline model for the robot agent constructed using our method is a text-based model, meaning the agent’s observations are input in textual form. However, there is a gap between textual descriptions of real-world visual images and the actual visual information, which cannot fully encapsulate all real-world details. This discrepancy affects the robot agent’s ability to ground itself in the environment. In future work, we aim to incorporate visual information directly into the input to better align with real-world scenarios. Additionally, although VirtualHome (Puig et al., 2018) is a relatively complex environment, we have not conducted experimental validation in other embodied environments or the real world. In the future, we will perform more experiments for validation.
@@
This work aims to construct a robot agent within Virtual Environment. The virtual environment setup and related data strictly follow the specifications of VirtualHome (Puig et al., 2018). We refer to VirtualHome v2.3.0111https://github.com/xavierpuigf/virtualhome/tree/master to conduct out our experiments (MIT license [^2] https://github.com/xavierpuigf/virtualhome/blob/master/LICENSE). The models, i.e. flan-t5-small, flan-t5-base and flan-t5-large (Chung et al., 2024), we use for fine-tuning are all open-source, and we will strictly follow the protocols for the academic use of these language models (Apache License 2.0333https://huggingface.co/google/flan-t5-base). In addition, we partially use AI assistants, such as Copilot and ChatGPT, to help with our coding and writing.
@@
Appendix A Data Template
We show the data template of planning data ($D_{p}$), feedback data ($D_{f}$) and correction data ($D_{c}$). Each piece of data adheres to the specified format.
Figure 7: Data template of planning data, feedback data and correction data
Appendix B Illustration of Error Type
Unflipped Boolean State error occurs when an action meant to change the state of an object with a Boolean attribute (such as open/closed or on/off) does not achieve the intended effect, like attempting to open an already open door. Missing Object error arises when the agent is not holding the necessary object to complete an action, preventing the task’s execution. Enclosed Object error involves the target object being contained within a closed structure, with the action failing to free the object for use. Invalid Action error occurs when the agent attempts to perform an action on a target object that is not afforded to it, such as trying to pull a ceiling. Over-occupied Agent error happens when the agent’s hands are occupied or already interacting with objects, leaving it unable to interact with the target object in the current step. Agent Proximity errors arise when the agent is not close enough to the target object to perform the action. Object availability errors occur when the agent attempts to interact with an object that does not exist in the environment. The remaining errors are categorized under Others.
Appendix C Length Analysis
Following our common sense, tasks with a greater number of steps are generally considered more challenging for the agent. To evaluate the performance of the agent on tasks of varying difficulty, we collected and analyzed the executability of tasks with different lengths of generated steps between our method and BC. As shown in  8, in terms of execution rates for different lengths of generated steps, our method outperforms BC, particularly in tasks with longer steps. This indicates the widespread efficacy of our method.
Appendix D More Details of Virtualhome Environment
VirtualHome provides diverse and customizable household environments that support a wide array of possible interactions in the form of atomic action steps. There are three kinds of action template based on the action type, which are "[Action]", "[Action] <Object> <id>" and "[Action] <Object> <id> <Object> <id>". Each [Action] refers to one of 42 atomic actions supported in Virtualhome. Full list of atomic actions are shown in  Table 5. In each scene, there are around 350 objects which the embodied agent can interact with and each object refers to the specific <id>. Each object has properties(e.g. drinkable, eatable) corresponding to its action affordances. Some objects also have a semantic state like heated, washed, or used.
In ActivityPrograms Knowledge Base (Puig et al., 2018), there are 292 unique high-level household tasks, with 1374 unique action plans and 6201 unique environments in total extracted from VirtualHome, and task and action plan samples manually annotated by Amazon Mechanical Turk workers. Each data sample consists of a high-level task, the description of the task and the complete action programs, which can be directly executed in the VirtualHome environment. A piece of data sample is shown in  Table 4.
@@ Appendix D More Details of Virtualhome Environment
Appendix E Further Experiment Details
@@
All experiments were conducted on eight NVIDIA RTX A6000 GPUs. During the pre-tuning phase, we selected 1000 samples from the expert planning data $D_{p}$ and trained for one epoch. During the training process, we set the following hyperparameters: a batch size of 30, training for three epochs, and selecting the best-performing checkpoints from these epochs. The learning rate was set to 1e-4. During the inference process, all generation parameters were kept consistent with the default generation parameters of the flan-t5 series models. All experiments are expected to reproduce in one day.
Input: $\mathcal{D}_{p}$: Expert planning data, $\pi_{\theta}$: initial robot agent policy, $T_{1}$: number of epochs in pre-tuning phase, VS: Virtualhome Simulator, $M$: Number of tasks, $n_{j}$: the step length of task j, M: GPT-4o, $T_{2}$: number of epochs in training phase.
@@
Randomly select few planning training data $\mathcal{D}_{few}\subseteq\mathcal{D}_{p}$
for $i=1$ to $T_{1}$ do
       Optimize $\theta$ on BC objective: $\mathcal{L}(\theta)=\mathbb{E}_{(q_{p},j_{t})\sim\mathcal{D}_{p}}\left[-\log\pi_{\theta}\left(a_{t}\mid(q_{p},j_{t-1})\right)\right]$
@@
for $j=1$ to $M$ do
       for $k=1$ to $n_{j}$ do
             Predicting the action $\hat{a}_{k}\sim\pi_{\theta}(q_{p},j_{k-1})$
             $\hat{a}_{k}$ executed in VS and obtain new observation $o_{k}$, environmental execution feedback $f_{k}$
@@
                   Correction data sample: ($q_{c},j_{k-1},\hat{a}_{k},f_{k},a_{k}$) added to $D_{c}$
                   Feedback data sample: ($q_{f},j_{k-1},\hat{a}_{k},f_{k}$) added to $D_{f}$
                   $a_{k}$ executed in VS and obtain new observation $o_{k}$
            if $\hat{a}_{k}==a_{k}$ then
                  Feedback data sample: ($q_{f},j_{k-1},a_{k},True$) added to $D_{f}$
@@
             Predicting the action $\hat{a}_{k}\sim\pi_{\theta}(q_{p},j_{k-1})$
             $\hat{a}_{k}$ executed in VS and obtain new observation $o_{k}$, environmental execution feedback $f_{k}$
             if $\hat{a}_{k}$ is non-executable then
                   Gain corrected action $a_{c}\sim M(q_{c},j_{k-1},\hat{a}_{k},f_{k})$
                   Correction data sample: ($q_{c},j_{k-1},\hat{a}_{k},f_{k},a_{c}$) added to $D_{c}$
            Feedback data sample: ($q_{f},j_{k-1},\hat{a}_{k},f_{k}$) added to $D_{f}$
// Learning from Exploration Past Experience
for $i=1$ to $T_{2}$ do
@@
Appendix F Pseudocode
This section presents the pseudocode of $E^{2}CL$ in Algorithm 2. A detailed discussion of the method is given in  Section 2.2.

### v2

@@ 3.3 Ablation Study on Training Data
In this section, we explore the impact of the collected training data, i.e., feedback data $D_{f}$ and correction data $D_{c}$, on overall performance by ablating them in the training. During the inference phase, we employ speculative inference for all settings to ensure consistency.
As shown in Table 2, we observe that both $D_{f}$ and $D_{c}$ are each beneficial for the agent, but lag behind the combination of them. We hypothesize that the improvement observed when training with $D_{c}$ is primarily due to the enhanced self-correction capability of the agent. However, the limited ability to generate high-quality action feedback hampers the effectiveness of self-correction during speculative inference, as demonstrated in Section 3.6. Compared to the agent training without both $D_{f}$ and $D_{c}$, training with $D_{p}$ and $D_{f}$ improves the performance by training to predict environmental feedback, which explicitly aligns with environmental knowledge. However, the weak self-correction capability of the agent constrains the agent from generating executable and correct action in speculative inference, which is demonstrated in  Section 3.5. In our method, we integrate both types of data, enabling our agent to generate higher-quality action feedback and exhibit stronger self-correction abilities. This results in a substantial performance boost compared to other ablation settings.
@@ 3.4 Analysis on Different Size of the Model
To investigate the impact of model size on performance, we train models of different sizes using both BC and our method, and evaluated them on unseen tasks. Our results are shown in Figure 3, where larger models perform relatively better across all aspects, indicating that model scale significantly impacts performance. Moreover, it can also be observed that our method outperforms BC in both affordance rate and LCS across models with different parameter sizes, which demonstrates that our method consistently provides superior performance regardless of model size. Notably, when using our method, smaller models achieve performance surpassing larger models using BC across all metrics. This finding suggests that our method is able to release the potential of small language models and lays the foundation for building agents that work on edge devices in the future.
@@ 3.5 Evaluation on Self-Correction Ability
Table 3: Performance of the agent regarding speculative inference (SI).
3.6 Analysis on Speculative Inference
To analyze the contribution of speculative inference to overall performance, as well as to explore the quality and effectiveness of self-generated feedback, we conduct the following analysis.
@@ 3.7 Error Analysis
We also perform an error analysis to identify the aspects where the agent constructed using our method outperforms BC. There are a total of eight types of errors, which can be further classified into grounding errors (object availability) and execution-related errors (others). The detailed demonstration can be found in the Appendix C.1. As shown in Figure 6, we observe that all error types decreased by more than 24%, with Over occupied error showing the highest reduction rate of 94.4%. This demonstrates the effectiveness of our method in reducing various types of errors, highlighting its comprehensiveness. For the two most frequent types of execution-related errors, unflipped boolean state and agent proximity, our method achieves a reduction in error count by over 37% compared to BC, thereby demonstrating its effectiveness. Although our method primarily aims to avoid execution errors related to physical constraint and does not specifically target grounding errors such as object availability, the fact that it still reduces this type of error demonstrates the generalizability of our method.
@@ 4 Related Work
After exploration, the agent would encounter failure in the past experience, which is assumed as negative samples. The topic of learning from negative samples has increasingly gained attention as an alternative approach to learning solely from positive samples. Traditionally, some studies aim to decrease the probability of negative samples while increasing the probability of positive samples in order to achieve better performance (Wang et al., 2024a; Zheng et al., 2023; Liu et al., 2022). Additionally, some works construct correcting dataset and tuning language models on these data (An et al., 2023; Wang et al., 2024b; Bai et al., 2022). Besides, there are other efforts aimed at leveraging the comprehension abilities of language models to widen the gap between positive and negative samples (Liu et al., 2023; Zhang et al., 2023; Tong et al., 2024). In our work, we similarly leverage the inherent understanding capabilities of language models and enhance the embodied agent’s learning from environmental feedback regarding exploration errors, as well as its ability to self-correct.
@@ 5 Conclusion
In this work, we aim to align the embodied agent with the environment to enhance its task-solving performance. Firstly, we present E2CL, a novel framework that leverages exploration-induced errors and environmental feedback to enhance environment alignment for LM-based agents during teacher-guided and teacher-free exploration. Furthermore, we introduce speculative inference, a process in which the agent utilizes learned abilities for self-feedback and self-correction to reduce execution errors. Extensive experiments show that our method outperforms many existing baseline methods.
@@
The baseline model for the embodied agent constructed using our method is a text-based model, meaning the agent’s observations are input in textual form. However, there is a gap between textual descriptions of real-world visual images and the actual visual information, which cannot fully encapsulate all real-world details. This discrepancy affects the robot agent’s ability to ground itself in the environment. In future work, we aim to incorporate visual information directly into the input to better align with real-world scenarios. Additionally, although VirtualHome (Puig et al., 2018) is a relatively complex environment, we have not conducted experimental validation in other embodied environments or the real world. In the future, we will perform more experiments for validation.
@@
This work aims to construct an embodied agent within Virtual Environment. The virtual environment setup and related data strictly follow the specifications of VirtualHome (Puig et al., 2018). We refer to VirtualHome v2.3.0111https://github.com/xavierpuigf/virtualhome/tree/master to conduct out our experiments (MIT license [^2] https://github.com/xavierpuigf/virtualhome/blob/master/LICENSE). The models, i.e. flan-t5-small, flan-t5-base and flan-t5-large (Chung et al., 2024), we use for fine-tuning are all open-source, and we will strictly follow the protocols for the academic use of these language models (Apache License 2.0333https://huggingface.co/google/flan-t5-base). In addition, we partially use AI assistants, such as Copilot and ChatGPT, to help with our coding and writing.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (62076212) and the Research Grants Council of Hong Kong (15209724). The authors would like to thank the anonymous reviewers for their valuable feedback and constructive suggestions.
@@ Acknowledgements
Appendix A Data
Figure 7: Data templates of planning data, feedback data and correction data.
Figure 8: Data examples of planning data, feedback data and correction data.
Appendix B Additional Details of VirtualHome Environment
VirtualHome provides diverse and customizable household environments that support a wide array of possible interactions in the form of atomic action steps. There are three kinds of action template based on the action type, which are "[Action]", "[Action] <Object> <id>" and "[Action] <Object> <id> <Object> <id>". Each [Action] refers to one of 42 atomic actions supported in Virtualhome. Full list of atomic actions are shown in  Table 6. In each scene, there are approximately 350 objects with which the embodied agent can interact, each identified by a specific <id>. These objects have properties (e.g., drinkable, eatable) corresponding to their action affordances. Some objects also possess a semantic state, such as heated, washed, or used.
In ActivityPrograms  (Puig et al., 2018) knowledge base, there are 292 unique high-level household tasks, with 1374 unique action plans and 6201 unique environments in total extracted from VirtualHome, and task and action plan samples manually annotated by Amazon Mechanical Turk workers. Each data sample consists of a high-level task, a description of the task, and complete action programs that can be directly executed in the VirtualHome environment. A piece of data sample is shown in  Table 4.
@@ Appendix B Additional Details of VirtualHome Environment
Appendix C Additional Analyses of Experiment Results
C.1 Illustration of Error Types
Unflipped Boolean State error occurs when an action meant to change the state of an object with a Boolean attribute (such as open/closed or on/off) does not achieve the intended effect, like attempting to open an already open door. Missing Object error arises when the agent is not holding the necessary object to complete an action, preventing the task’s execution. Enclosed Object error occurs when the target object is contained within a closed structure, preventing the action from freeing the object for use. Invalid Action error occurs when the agent attempts to perform an action on a target object that is not afforded to it, such as trying to pull a ceiling. Over-occupied Agent error happens when the agent’s hands are occupied or already interacting with objects, leaving it unable to interact with the target object in the current step. Agent Proximity errors arise when the agent is not close enough to the target object to perform the action. Object availability errors occur when the agent attempts to interact with an object that does not exist in the environment. The remaining errors are categorized as Others.
C.2 Length Analysis
Following our common sense, tasks with a greater number of steps are generally considered more challenging for the agent. To evaluate the performance of the agent on tasks of varying difficulty, we collected and analyzed the executability of tasks with different lengths of generated steps between our method and BC. As shown in Figure 9, in terms of execution rates for different lengths of generated steps, our method outperforms BC, particularly in tasks with longer steps. This indicates the widespread efficacy of our method.
C.3 Convergence Speed Analysis
We report the variations of training losses among our method and two representative baseline methods. The results are shown in the Table 5. We observe that as the number of epochs increases, the loss for our method decreases more rapidly than that of BC and NAT, ultimately converging at a lower bound. It indicates that our method contributes to the convergence speed during model training.
Loss  epoch 1  epoch 2  epoch 3  epoch 4  epoch 5  epoch 6  epoch 7
BC 4.4470 0.2956 0.2631 0.1987 0.1476 0.1075 0.0967
NAT 5.8337 0.2164 0.1308 0.1207 0.0758 0.0600 0.0503
Ours 5.3267 0.1252 0.1134 0.0849 0.0708 0.0423 0.0354
Table 5: Comparison of convergence speed between our method and baseline methods
@@ C.3 Convergence Speed Analysis
Appendix D Additional Implementation Details
@@
All experiments were conducted on eight NVIDIA RTX A6000 GPUs. During the pre-tuning phase, we selected 1000 samples from the expert planning data $D_{p}$ and trained for one epoch. During the training process, we set the following hyperparameters: a batch size of 30, training for three epochs, and selecting the best-performing checkpoints from these epochs. The learning rate was set to 1e-4. During the inference process, all generation parameters were kept consistent with the default generation parameters of the flan-t5 series models. The results reported in the paper are all averages. All experiments are expected to reproduce in one day.
Input: $\mathcal{D}_{p}$: Expert planning data, $\pi_{\theta}$: initial robot agent policy, $T_{1}$: number of epochs in pre-tuning phase, ES: Environment Simulator, $|\mathcal{J}|$: Number of tasks, $n_{i}$: the step length of task i, M: GPT-4o, $T_{2}$: number of epochs in training phase.
@@
Randomly select few planning training data $\mathcal{D}_{p^{\prime}}\subseteq\mathcal{D}_{p}$
for $h=1$ to $T_{1}$ do
       Optimize $\theta$ on BC objective: $\mathcal{L}(\theta)=\mathbb{E}_{(q_{p},j_{t})\sim\mathcal{D}_{p^{\prime}}}\left[-\log\pi_{\theta}\left(a_{t}\mid(q_{p},j_{t-1})\right)\right]$
@@
for $i=1$ to $|\mathcal{J}|$ do
       for $t=1$ to $n_{i}$ do
             Predicting the action $\hat{a}_{t}\sim\pi_{\theta}(q_{p},j_{t-1})$
             $\hat{a}_{t}$ executed in ES and obtain new observation $o_{t}$, environmental execution feedback $f_{t}$
@@
                   Correction data sample: ($q_{c},j_{t-1},\hat{a}_{t},f_{t},a_{t}$) added to $D_{c}$
                   Feedback data sample: ($q_{f},j_{t-1},\hat{a}_{t},f_{t}$) added to $D_{f}$
                   $a_{t}$ executed in ES and obtain new observation $o_{t}$
            if $\hat{a}_{t}==a_{t}$ then
                  Feedback data sample: ($q_{f},j_{t-1},a_{t},True$) added to $D_{f}$
@@
             Predicting the action $\hat{a}_{t}\sim\pi_{\theta}(q_{p},j_{t-1})$
             $\hat{a}_{t}$ executed in ES and obtain new observation $o_{t}$, environmental execution feedback $f_{t}$
             if $\hat{a}_{t}$ is non-executable then
                   Gain corrected action $a_{t}\sim M(q_{c},j_{t-1},\hat{a}_{t},f_{t})$
                   Correction data sample: ($q_{c},j_{t-1},\hat{a}_{t},f_{t},a_{t}$) added to $D_{c}$
            Feedback data sample: ($q_{f},j_{t-1},\hat{a}_{t},f_{t}$) added to $D_{f}$
// Learning from Exploration Experience
for $h=1$ to $T_{2}$ do
@@
Appendix E Pseudocode
This section presents the pseudocode of E2CL in  Algorithm 2. A detailed discussion of the method is given in  Section 2.2.